# Metric-Based Few-Shot Learning for Exercise Repetition Counting with IMU Data

### v1

@@ Abstract
IMU sensing technology has significantly advanced in automating systems that reflect human movement, particularly in robotics and healthcare, where it is used to automatically detect target movements. This study develops a method to automatically count exercise repetitions by analyzing IMU signals, with a focus on a universal exercise repetition counting task that counts all types of exercise movements, including novel exercises not seen during training, using a single model. A key challenge in developing such a model is handling the considerable variation in peak patterns across different types of exercises. Since peak patterns can vary significantly between different exercises as well as between individuals performing the same exercise, the model needs to learn a complex embedding space of sensor data to generalize effectively.
To address this challenge, we propose a repetition counting technique utilizing a deep metric-based few-shot learning approach, designed to handle both existing and novel exercises. By redefining the counting task as a few-shot classification problem, the method is capable of detecting peak repetition patterns in exercises not seen during training. The approach employs a Siamese network with triplet loss, optimizing the embedding space to distinguish between peak and non-peak frames. The proposed framework is composed of three main phases: standard classification training, few-shot training, and fine-tuning for novel exercises, followed by post-processing to refine the final repetition counts. Evaluation results demonstrate the effectiveness of the proposed approach, showing an 86.8% probability of accurately counting ten or more repetitions within a single set across 28 different exercises. This performance highlights the model’s ability to generalize across various exercise types, including those not present in the training data. Such robustness and adaptability make the system a strong candidate for real-time implementation in fitness and healthcare applications
{IEEEkeywords}
Few-Shot Learning, Human Activity Recognition, Metric-based Learning, Repetition Counting, IMU Sensor
@@ 4.3 Few-Shot Learning for Repetition Counting
{IEEEeqnarray}
C ^y_i = Sigmoid(W_o~x_i + b_o)
L_phase1=-1N∑_i=1^N y_i log(^y_i)+(1-y_i)log(1-^y_i)
@@
To avoid inefficiency in using all possible triplet combinations, we apply semi-hard triplet mining [15]. This method selects the negative sample ($x^{n}$) such that it is not closer to the anchor ($x^{a}$) than the positive sample ($x^{p}$), and the distance between $x^{a}$ and $x^{n}$ does not exceed the distance between $x^{a}$ and $x^{p}$ by more than a margin ($\alpha$), as shown in equation (2). This ensures that extremely easy or difficult samples are excluded, focusing the training on moderately challenging examples. The margin $\alpha$ is set to 1.0.
$\begin{array}[]{cl}\small\mathcal{}{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}<{\|f_{\theta}(x^{a})-f_{\theta}(x^{n})\|}^{2}\\ <{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}+\alpha\end{array}$ (2)
The triplet loss, defined in equation (3), aims to bring the positive pair closer in the embedding space and push the negative pair farther apart. The hinge loss function $[\cdot]_{+}$ ensures that only the violations of the margin contribute to the loss, meaning that when the negative sample is already sufficiently far from the anchor, no additional penalty is applied.
$\small\mathcal{}L_{triplet}(t)=[{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}-{\|f_{\theta}(x^{a})-f_{\theta}(x^{n})\|}^{2}+\alpha]_{+}$ (3)
The total triplet loss across all triplet pairs is averaged, as shown in equation (4), to optimize the model’s performance on the dataset.
$\small\mathcal{}L_{phase2}=\frac{1}{N}\sum\limits_{i=1}^{N}L_{triplet}(t_{i})$ (4)
@@
Post-processing In this section, we discuss how classification results are determined and repetitions are counted. Using the $f_{\theta}$ obtained from Phase 3, we compare the similarities between two pairs, $(x^{a},x^{p})$ and $(x^{a},x^{n})$, and classify $x^{a}$ into the class with the higher similarity. The cosine similarity of the positive pair ($SP$) and the negative pair ($SN$) is calculated using equation (5) to determine peak and non-peak points. $x^{p}$ and $x^{n}$ are randomly selected from the set constructed from $D_{register}$, and the final similarity is averaged over five samples. The class is then determined by comparing $SP$ and $SN$.
$\begin{array}[]{cl}\small\mathcal{}SP=\frac{x^{a}\cdot x^{p}}{\|x^{a}\|\|x^{p}\|}\quad SN=\frac{x^{a}\cdot x^{n}}{\|x^{a}\|\|x^{n}\|}\\ \\ \hat{y}=\begin{cases}\textrm{peak},\;\textrm{if}\;SP\geq SN\\ \textrm{non-peak},\;\textrm{if}\;SP<SN\end{cases}\par\end{array}$ (5)
@@ 5.2 Results on Repetition Counting
Table 4 presents the results of repetition counting for individual exercises in terms of the number of errors compared to the actual count using the Siamese network. The columns indicate the repetition error ratio. “e$|X|$” denotes the number of exercise sets with the “$|X|$” repetition error count. “$|X|$” represents the absolute error count of 0, 1, 2, 3, 4, 5, or more than five errors. In our dataset, the mean repetitions per one set is fifteen. The percentage of errors more than five times is 13.2%, and the rate of error-free counting is 12.9%. The probability of correctly calculating ten or more repetitions in one set is 86.8%. Generally, nine repetitions in one set were counted accurately.

### v2

@@ Abstract
Sensing technology has significantly advanced in automating systems that reflect human movement, particularly in robotics and healthcare, where it is used to automatically detect target movements. This study develops a method to automatically count exercise repetitions by analyzing IMU signals, with a focus on a universal exercise repetition counting task that counts all types of exercise movements, including novel exercises not seen during training, using a single model. Since peak patterns can vary significantly between different exercises as well as between individuals performing the same exercise, the model needs to learn a complex embedding space of sensor data to generalize effectively.
To address this challenge, we propose a repetition counting technique utilizing a deep metric-based few-shot learning approach, designed to handle both existing and novel exercises. By redefining the counting task as a few-shot classification problem, the method is capable of detecting peak repetition patterns in exercises not seen during training. The approach employs a Siamese network with triplet loss, optimizing the embedding space to distinguish between peak and non-peak frames. Evaluation results demonstrate the effectiveness of the proposed approach, showing an 86.8% probability of accurately counting ten or more repetitions within a single set across 28 different exercises. This performance highlights the model’s ability to generalize across various exercise types, including those not present in the training data. Such robustness and adaptability make the system a strong candidate for real-time implementation in fitness and healthcare applications
keywords:
[…]
@@ 4.3 Few-shot learning for repetition counting
$\begin{array}[]{cl}\hat{y}_{i}=Sigmoid(W_{o}\tilde{x}_{i}+b_{o})\\ L_{phase1}=-\frac{1}{N}\sum\limits_{i=1}^{N}y_{i}log(\hat{y}_{i})+(1-y_{i})log(1-\hat{y}_{i})\end{array}$ (2)
@@
To avoid inefficiency in using all possible triplet combinations, we apply semi-hard triplet mining [19]. This method selects the negative sample ($x^{n}$) such that it is not closer to the anchor ($x^{a}$) than the positive sample ($x^{p}$), and the distance between $x^{a}$ and $x^{n}$ does not exceed the distance between $x^{a}$ and $x^{p}$ by more than a margin ($\alpha$), as shown in Eq. (3). This ensures that extremely easy or difficult samples are excluded, focusing the training on moderately challenging examples. The margin $\alpha$ is set to 1.0.
$\begin{array}[]{cl}{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}<{\|f_{\theta}(x^{a})-f_{\theta}(x^{n})\|}^{2}\\ <{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}+\alpha\end{array}$ (3)
The triplet loss, defined in Eq. (4), aims to bring the positive pair closer in the embedding space and push the negative pair farther apart. The hinge loss function $[\cdot]_{+}$ ensures that only the violations of the margin contribute to the loss, meaning that when the negative sample is already sufficiently far from the anchor, no additional penalty is applied.
$L_{triplet}(t)=[{\|f_{\theta}(x^{a})-f_{\theta}(x^{p})\|}^{2}-{\|f_{\theta}(x^{a})-f_{\theta}(x^{n})\|}^{2}+\alpha]_{+}$ (4)
The total triplet loss across all triplet pairs is averaged, as shown in Eq. (5), to optimize the model’s performance on the dataset.
$L_{phase2}=\frac{1}{N}\sum\limits_{i=1}^{N}L_{triplet}(t_{i})$ (5)
@@
Post-processing In this section, we discuss how classification results are determined and repetitions are counted. Using the $f_{\theta}$ obtained from Phase 3, we compare the similarities between two pairs, $(x^{a},x^{p})$ and $(x^{a},x^{n})$, and classify $x^{a}$ into the class with the higher similarity. The cosine similarity of the positive pair ($SP$) and the negative pair ($SN$) is calculated using Eq. (6) to determine peak and non-peak points. $x^{p}$ and $x^{n}$ are randomly selected from the set constructed from $D_{register}$, and the final similarity is averaged over five samples. The class is then determined by comparing $SP$ and $SN$.
$\begin{array}[]{cl}SP=\frac{x^{a}\cdot x^{p}}{\|x^{a}\|\|x^{p}\|}\quad SN=\frac{x^{a}\cdot x^{n}}{\|x^{a}\|\|x^{n}\|}\\ \\ \hat{y}=\begin{cases}\textrm{peak},\;\textrm{if}\;SP\geq SN\\ \textrm{non-peak},\;\textrm{if}\;SP<SN\end{cases}\par\end{array}$ (6)
@@ 5.2 Results on repetition counting
Table 4 presents the results of repetition counting for individual exercises in terms of the number of errors compared to the actual count using the Siamese network. The columns indicate the repetition error ratio. “e$|X|$" denotes the number of exercise sets with the “$|X|$” repetition error count. “$|X|$” represents the absolute error count of 0, 1, 2, 3, 4, 5, or more than five errors. In our dataset, the mean repetitions per one set is fifteen. The percentage of errors more than five times is 13.2%, and the rate of error-free counting is 12.9%. The probability of correctly calculating ten or more repetitions in one set is 86.8%. Generally, nine repetitions in one set were counted accurately.